# Exceptional-point-controlled mode interaction in three-dimensional microcavities represented by generalized Husimi functions

### v1

@@ III.4 Wave and Husimi function morphology evolution in conical cylinders
{tblr}
c—[dashed]c—[dashed]c & $\delta=0$ $\delta=0.1$
[…]
Figure 5: $E_{\parallel}$ field intensity (left) and Husimi functions at $z=0$ (right) of TM modes with (a,b) $\eta=0$, (c,d) $\eta=3$ and (e,f) $\eta=7$ for deformation parameters (a,c,e) $\delta=0$ and (b,d,f) $\delta=0.1$.
@@
with deformation strength $\delta=\Delta R/R_{0}$ and $z\in[0,h]$. In the following, we focus on systems with $h/R_{0}=5$ and modes with azimuthal quantum number $m=6$. This geometry can also be described as a cone stump with opening angle $\alpha$ where $\tan\alpha=\Delta R/h=\delta/5$, see Fig. III.4a. This geometry is relevant for experimental setups, as microcavity side walls are usually slightly slanted due to fabrication methods. Fig. III.4b shows the resulting effective optical potential
@@
The left panels in Fig. III.4c-h show the field functions of three different TM modes with varying excitations in the $z$-direction. The real parts of the eigenfrequencies decrease for all modes. This is due to $n\mathrm{Re}\,(\Omega)R_{\mathrm{eff}}/R_{0}\approx\mathrm{const}$ for a given mode. The deformation $\delta$ increases the effective radius, with $R_{0}(1+\delta)\propto R_{\mathrm{eff}}$, which, in turn, implies $1+\delta\propto\mathrm{Re}\,(\Omega)^{-1}$. The field intensity shifts to the bottom of the cavity for low $\eta$ (see Figs. III.4c,d and e,f) due to the lower optical potential according to Eq. (12). For modes with high $\eta$, i.e. Figs. III.4g and h, the morphological adaptability is constrained by the necessity to remain orthogonal to the modes with lower $\eta$. The energetically lower states already occupy the space at the bottom where of the radius is largest, making it less accessible for the higher states.
@@
In Fig. III.4f and h, we see that the field concentrates in the top intensity maximum; this is also a byproduct of the orthogonality requirement. Hence, less field intensity leaks out of the cavity top face for modes with low $\eta$, as the field intensity is very low in that area. The imaginary component of $\Omega$ decreases accordingly. In contrast, modes with high $\eta$ exhibit an increase of the imaginary component. The proximity of the uppermost intensity maximum to the cylinder top face results in a reduced mode lifetime due to increased evanescent losses.
@@
Figure 6: Real parts of eigenvalues for the slanted cone. All modes with $\mathrm{Re}\,\Omega\in[4.5,5.5]$ are shown. Color denotes polarization $P\in[-1,1]$ according to Eq. 14, with red (blue) corresponding to TM (TE) polarized modes. Grey dots show unpolarized modes with $P\approx 0$. The spectrum at $\delta=0$ reflects the undeformed cylinder. The dark grey area marks the location of the avoided resonance crossing in Fig. 7.
Husimi functions at $z=h/2$ are shown in the right hand panels of Fig. III.4c-h. Here, the wave function on the cavity surface was projected onto a Euclidean plane, as discussed in Sec. III.2. Similarly to the cylinder, the cone surface can be unrolled using an appropriate transformation, $\psi(\varphi,z)\rightarrow\psi(s,\tilde{z})$, where $\tilde{z}$ is the coordinate in the unrolled plane. The details can be found in Appendix B.
@@
For $\delta=0$ in Figs. III.4c,e and g, the behavior corresponds to $\eta$ as discussed above. Due to the mirror symmetry at $z=h/2$, the phase space is symmetric with respect to $\sin\chi_{z}=0$. Generally, this changes for nonzero $\delta$. The broken symmetry is visible in Figs. III.4d,f and h, where the Husimi functions for $\delta=0.1$ are shown. The signature of the mode with $\eta=0$ in Fig. III.4d is only slightly distorted in comparison to Fig. III.4c. Even with the change in morphology, the predominant motion is still along the cavity boundary, resulting in only the signature at $\sin\chi_{z}=0$. In Fig. III.4f the asymmetry is more pronounced for the mode with $\eta=2$. The uppermost field-intensity maximum is dominant in size and intensity due to the deformation. Due to this spatial expansion, the two Husimi maxima merge into one, which is shifted into the area of $\sin\chi_{z}<0$. The overall intensity inside the cavity at $z>h/2$ is drastically lower than for $z<h/2$ due to the mode shift into areas of higher radius. For high $\eta$, e.g. $\eta=7$ in Fig. III.4h, the phase space distribution is more balanced, which is a result of the more evenly distributed field function. However, a distortion of the phase space in comparison to Fig. III.4g is visible, emphasizing that the mode structure is more complex for the deformed cavity. It is evident that the Husimi function encodes the quantum numbers and the local morphology of a mode.
@@ IV.1 TE-TM mode coupling and formation of exceptional points
(c) $\mathrm{Re}\,\Omega$
(d) $\mathrm{Im}\,\Omega$
Figure 7: Eigenvalue evolution of $(2,6)$-TM and $(2,2)$-TE modes. The (a) real and (b) imaginary components of the eigenfrequency, $\Omega$. The annotated letters relate to Fig. IV.1 (IV.2) for the field (Husimi) functions at the respective parameters. (c,d) Riemann surface around the EP, which governs the coupling between the two modes for (c) $\mathrm{Re}\,\Omega$ and (d) $\mathrm{Im}\,\Omega$. The gray dots correspond to data used to calculate the energy landscape. The black dots denote the positions of the EPs, which are located at ${\delta\approx 0.2217}$ and ${n\approx 2.4335}$. The black solid lines surrounding the data points illustrate the non-trivial shapes of the surfaces. The black dashed lines denote the branch cuts with $\mathrm{Re}\,\Omega\,(\mathrm{Im}\,\Omega)=0$.
@@
$\mathcal{H}=\begin{pmatrix}\Omega^{(1)}&V\\ W&\Omega^{(2)}\end{pmatrix},\quad$ (17)
where $\Omega^{(1,2)}$ are the frequencies of the uncoupled modes and $V,W$ are the coupling coefficients. The two modes can only interact meaningfully if $\sqrt{|VW|}\approx|\Omega^{(1)}-\Omega^{(2)}|$. If this condition is satisfied, an avoided resonance crossing (ARC) occurs in either the real or imaginary component of $\Omega$, implying a level crossing in the other component. An exceptional point (EP) is present when $V=0$ or $W=0$, which is accompanied by a coalescence of the two resulting eigenvalues and vectors.[67, 68]
@@
One question to explore is why TM-TE ARCs are the most common type. First, let us consider the factors that exclude ARCs from occurring between two modes. Initially, we focus on TM-TM and TE-TE modes with identical radial quantum number, $\rho$. Modes with lower $\eta$ tend to concentrate more of the mode in the region of larger radius (see Fig. III.4). This leads to monotonically increasing $\mathrm{Re}\,\Omega(\eta)$ for a fixed $\delta$ in Fig. 6. This is the case for any $\delta$, hence no crossings can occur. As a consequence, only modes with different $\rho$ can exhibit crossings in the first place. Thus, different $\eta$ are required in order for two resonances to have similar values of $\mathrm{Re}\,\Omega$. This affects the mode lifetimes $\mathrm{Im}\,\Omega$, which must also be similar in order to minimize $|\Omega^{(1)}-\Omega^{(2)}|$. Due to these two aspects, TM-TM and TE-TE ARCs are very rare in comparison to TM-TE ARCs. The boundary conditions in Eqs. (5) lead to $\delta$ affecting the frequencies of TM/TE modes differently, thereby creating the possibility for mode interaction.
{tblr}
c—[dashed]c—[dashed]c upper branch lower branch
[dashed] $\delta=0$ Refer to caption Refer to caption
[dashed] $\delta=0.21$ Refer to caption Refer to caption
[dashed] $\delta=0.23$ Refer to caption Refer to caption
[dashed] $\delta=0.25$ Refer to caption Refer to caption
Figure 8: Field functions of (2,6)-TM and (2,2)-TE modes for (a,b) $\delta=0$ and (c-h) at the ARC (see Fig. 7a,b). The fields are normalized so that $\max|E_{j}|^{2}=1$ and $\max|B_{j}|^{2}=1$, where the subscript $j$ denotes the three field components.
In the following, we concentrate on a specific ARC between a $(1,6)$-TM mode and a $(1,2)$-TE mode, as it is the most common ARC type. Another ARC, where one TM and two TE modes are involved, is shown in the supplementary material in Appendix D. The eigenvalue evolution of the two modes close to the ARC is shown in Fig. 7a-b. First consider $\mathrm{Re}\,\Omega$ in Fig. 7a. The two frequency branches approach each other with increasing $\delta$. They are closest at $\delta\approx 0.23$, where both modes exhibit TE polarization. For $\delta<0.23$ the upper (lower) branch is TE (TM) polarized. They switch polarization for $\delta>0.23$. The imaginary component in Fig. IV.1b shows a level crossing. This is expected according to the effective Hamiltonian in Eq. 17. When an (anti-)crossing in (real) imaginary components is present, the coupling is called strong coupling. A second crossing also occurs in the imaginary part at $\delta\approx 0.255$ that is not related to the ARC being considered, but another ARC that is adjacent.
@@
The landscape, i.e. the Riemann surface, close to the EP of the $(1,6)$-TM and $(1,2)$-TE modes is shown in Fig. 7c-d. The solid black outlines show that the EP needs to be encircled twice in order to return to the starting position. The dashed lines show where $\mathrm{Re}\,(\mathrm{Im}\,)\Delta\Omega=0$, indicating the switch from strong ($\Delta\mathrm{Im}\,\Omega=0$) to weak coupling ($\Delta\mathrm{Re}\,\Omega=0$) when decreasing the refractive index. [67, 71] In the strong coupling regime the parameters in Eq. 17 are such that $2\sqrt{|VW|}>|\mathrm{Im}\,(\Omega^{(1)}-\Omega^{(2)})|$, which results in the level crossing occurring in $\mathrm{Re}\,\Omega$. [67, 71] This is the case for $n>2.4335$ in (see Figs. 7a,c) for the considered TM-TE mode pair. The weak coupling regime with $2\sqrt{|VW|}<|\mathrm{Im}\,(\Omega^{(1)}-\Omega^{(2)})|$ is present for $n<2.4335$ (see Fig. 7d).
The polarization changes are visible in the wave functions in Fig. IV.1. The upper branch for $\delta=0$ in Fig. IV.1a is distinctly TE polarized, as the out-of-plane component of the magnetic field dominates the mode structure. The lower branch is TM polarized, see Fig. IV.1b. When $\delta$ is increased, the change in mode morphology takes place as discussed in Sec. III.4. The intensity distribution of the TE mode shifts to lower $z$ where the radius increases due to its low $\eta$ (see Fig. IV.1c). The TM mode has an intensity maximum close to the top due to its high $\eta$ (see Fig. IV.1d). Near the ARC in Figs. IV.1e and f both $E_{\parallel}$ and $B_{\parallel}$ are of significant intensity. The TM and TE modes hybridize, which causes both TM and TE components being present. For larger deformations (see Figs. IV.1g,h) the polarization of the two branches switches.
It is evident that EPs govern the coupling between TM and TE modes. The mode pairs whose crossings do not show ARC behavior in Fig. 6 could perhaps be coupled if an appropriate deformation were chosen. Their coupling is also governed by EPs, even though their signatures are not visible here.
@@
{tblr}
c—[dashed]c—[dashed]c upper branch lower branch
[dashed] $\delta=0$ Refer to caption Refer to caption
[dashed] $\delta=0.21$ Refer to caption Refer to caption
[dashed] $\delta=0.23$ Refer to caption Refer to caption
[dashed] $\delta=0.25$ Refer to caption Refer to caption
Figure 9: Husimi functions of modes corresponding to ARC (see Figs. 7a,b). The red (blue) figures correspond to the $E_{\parallel}$ ($B_{\parallel}$) field component used for calculation The field functions are shown in Fig. IV.1, where the numerals in curly brackets refer to the respective parameter.
The Husimi function signatures of the coupled modes are shown in Fig. IV.2. We choose the Husimi function at $\bm{q}/h=(0,2.5)$, as nearly all wave functions have a nonzero intensity in the vicinity. The red (blue) colors show the Husimi functions calculated using $E_{\parallel}$ ($B_{\parallel}$). The signatures for the dominant field component of the uncoupled modes at $\delta=0$ are as expected from Fig. III.4.
The dominant out-of-plane component of the TE mode in Fig. IV.1a is the $B_{\parallel}$ field. Its Husimi function signature (see Fig. IV.2b) consists of two intensity maxima, which merge to form one broad intensity maximum. This corresponds to Fig. III.4c due to the low $\eta$ value. The low-intensity $E_{\parallel}$ field component is oscillating more in $z$, which results in a Husimi function structure resembling a resonance with high $\eta$ (compare to Fig. 4c).
The uneven spacing of the intensity maxima in $E_{\parallel}$ leads to broader peaks in the Husimi function. The phase space of the TM polarized mode in Fig. IV.1b (left panel) shows the typical behavior of $\eta=6$ in $E_{\parallel}$, as it corresponds to the behavior in Figs. 4c. Due to the small intensity fluctuations in the $B_{\parallel}$ component (right panel) the Husimi function yields four intensity maxima.
In Figs. IV.2c and d, the phase space signatures and the mode morphologies change according to Fig. III.4. Furthermore, both the $E_{\parallel}$ and $B_{\parallel}$ components become significant near the ARC. This is seen, for example, in Figs. IV.1e,f and IV.2e,f: The electric fields and phase space representations of both eigenmodes display TM mode characteristics, while the magnetic field (and its Husimi function) show TE behavior. The mode hybridization caused by the coupling is thus expressed in real and in phase space. When the system moves away from the ARC, the polarization of the modes increases. The Husimi function signatures in Figs. IV.2g,h reflect the behavior in Figs. IV.2c,d. Now the polarization is switched, which is reflected in phase space. For fields and Husimi functions at the EP we refer to Appendix E.

### v2

@@ III.4 Wave and Husimi function morphology evolution in conical cylinders
$\delta=0$ $\delta=0.1$
[dashed]
$\eta=0$
Refer to caption Ω= 4.820 - i  0.038 Refer to caption Ω= 4.487 - i  0.032
[dashed]
$\eta=2$
Refer to caption Ω= 4.854 - i  0.035 Refer to caption Ω= 4.630 - i  0.028
[dashed]
$\eta=7$
Refer to caption Ω= 5.185 - i  0.013 Refer to caption Ω= 5.002 - i  0.015
Figure 6: (a) System geometry and (b) optical potential of the slanted cone. (c-h) $E_{\parallel}$ field intensity (left) and Husimi functions at $z=0$ (right) of TM modes with (c,d) $\eta=0$, (e,f) $\eta=3$ and (g,h) $\eta=7$ for deformation parameters (c,e,g) $\delta=0$ and (d,f,h) $\delta=0.1$.
@@
with deformation strength $\delta=\Delta R/R_{0}$ and $z\in[0,h]$. In the following, we focus on systems with $h/R_{0}=5$ and modes with azimuthal quantum number $m=6$. This geometry can also be described as a cone stump with opening angle $\alpha$ where $\tan\alpha=\Delta R/h=\delta/5$, see Fig. 6(a). This geometry is relevant for experimental setups, as microcavity side walls are usually slightly slanted due to fabrication methods. Fig. 6(b) shows the resulting effective optical potential
@@
The left panels in Fig. 6(c-h) show the field functions of three different TM modes with varying excitations in the $z$-direction. The real parts of the eigenfrequencies decrease for all modes. This is due to $n\mathrm{Re}\,(\Omega)R_{\mathrm{eff}}/R_{0}\approx\mathrm{const}$ for any given mode. The deformation $\delta$ increases the effective radius, with $R_{0}(1+\delta)\propto R_{\mathrm{eff}}$, which, in turn, implies $1+\delta\propto\mathrm{Re}\,(\Omega)^{-1}$. The field intensity shifts to the bottom of the cavity for low $\eta$ (see Figs. 6(c,d) and (e,f)) due to the lower optical potential according to Eq. (12). For modes with high $\eta$, i.e. Figs. 6(g,h), the morphological adaptability is constrained by the necessity to remain orthogonal to the modes with lower $\eta$. The energetically lower states already occupy the space at the bottom where of the radius is largest, making it less accessible for the higher states.
@@
In Figs. 6(f,h), we see that the field concentrates in the top intensity maximum when $\delta$ is increased; this is also a byproduct of the orthogonality requirement. Hence, less field intensity leaks out of the cavity top face for modes with low $\eta$, as the field intensity is very low in that area. The imaginary component of $\Omega$ increases accordingly. In contrast, modes with high $\eta$ exhibit decreasing imaginary components. The proximity of the uppermost intensity maximum to the cylinder top face results in
increased losses, induced by the sharp corner. Similar behavior is also prevalent in 2D cavities, where sharp edges are major sources of refractive escape [70, 71]. With this mechanism, the deformation $\delta$ can in- or decrease the resonance lifetime, depending on the excitation $\eta$ in $z$.
@@
Figure 7: Real parts of eigenvalues for the slanted cone. All modes with $\mathrm{Re}\,\Omega\in[4.5,5.5]$ are shown. Color denotes polarization $P\in[-1,1]$ according to Eq. (14), with red (blue) corresponding to TM (TE) polarized modes. Grey dots show unpolarized modes with $P\approx 0$. The spectrum at $\delta=0$ reflects the undeformed cylinder. The dark grey area marks the location of the avoided resonance crossing in Fig. 8.
[…]
@@
For $\delta=0$ in Figs. 6(c,e,g), the behavior corresponds to $\eta$ as discussed above. Due to the mirror symmetry at $z=h/2$, the phase space is symmetric with respect to $\sin\chi_{z}=0$. Generally, this changes for nonzero $\delta$. The broken symmetry is visible in Figs. 6(d,f,h), where the Husimi functions for $\delta=0.1$ are shown. The signature of the mode with $\eta=0$ in Fig. 6(d) is only slightly distorted in comparison to Fig. 6(c). Even with the change in morphology, the predominant motion is still along the cavity boundary, resulting in only the signature at $\sin\chi_{z}=0$. In Fig. 6(f) the asymmetry is more pronounced for the mode with $\eta=2$. The uppermost field-intensity maximum is dominant in size and intensity due to the deformation. Due to this spatial expansion, the two Husimi maxima merge into one, which is shifted into the area of $\sin\chi_{z}<0$. The overall intensity inside the cavity at $z>h/2$ is drastically lower than for $z<h/2$ due to the mode shift into areas of higher radius. For high $\eta$, e.g. $\eta=7$ in Fig. 6(h), the phase space distribution is more balanced, which is a result of the more evenly distributed field function. However, a distortion of the phase space in comparison to Fig. 6(g) is visible, emphasizing that the mode structure is more complex for the deformed cavity. It is evident that the Husimi function encodes the quantum numbers and the local morphology of a mode.
@@ IV.1 TE-TM mode coupling and formation of exceptional points
Figure 8: Eigenvalue evolution of $(2,6)$-TM and $(2,2)$-TE modes. The (a) real and (b) imaginary components of the eigenfrequency, $\Omega$. The annotated letters relate to Fig. 9 (10) for the field (Husimi) functions at the respective parameters. (c,d) Riemann surface around the EP, which governs the coupling between the two modes for (c) $\mathrm{Re}\,\Omega$ and (d) $\mathrm{Im}\,\Omega$. The gray dots correspond to data used to calculate the energy landscape. The black dots denote the positions of the EP, which is located at $\delta\approx 0.2217$} and $n\approx 2.4335$}. The black solid lines surrounding the data points illustrate the non-trivial shapes of the surfaces. The black dashed lines denote the branch cuts with $\mathrm{Re}\,\Omega\,(\mathrm{Im}\,\Omega)=0$.
@@
$\mathcal{H}=\begin{pmatrix}\Omega^{(1)}&V\\ W&\Omega^{(2)}\end{pmatrix}$$\quad$ (17)
where $\Omega^{(1,2)}$ are the frequencies of the uncoupled modes and $V,W$ are the coupling coefficients. The two modes can only interact meaningfully if $\sqrt{|VW|}\approx|\Omega^{(1)}-\Omega^{(2)}|$. If this condition is satisfied, an avoided resonance crossing (ARC) occurs in either the real or imaginary component of $\Omega$, implying a level crossing in the other component. An exceptional point (EP) is present when $V=0$ or $W=0$, which is accompanied by a coalescence of the two resulting eigenvalues and vectors [72, 73].
@@
One question to explore is why TM-TE ARCs are the most common type. First, let us consider the factors that exclude ARCs from occurring between two modes. Initially, we focus on TM-TM and TE-TE modes with identical radial quantum number, $\rho$. Modes with lower $\eta$ tend to concentrate more of the mode in the region of larger radius (see Fig. 6). This leads to monotonically increasing $\mathrm{Re}\,\Omega(\eta)$ for a fixed $\delta$ in Fig. 7. This is the case for any $\delta$, hence no crossings can occur. As a consequence, only modes with different $\rho$ can exhibit crossings in the first place. Thus, different $\eta$ are required in order for two resonances to have similar values of $\mathrm{Re}\,\Omega$. This affects the mode lifetimes $\mathrm{Im}\,\Omega$, which must also be similar in order to minimize $|\Omega^{(1)}-\Omega^{(2)}|$. Due to these two aspects, TM-TM and TE-TE ARCs are very rare in comparison to TM-TE ARCs. The boundary conditions in Eqs. (5) lead to $\delta$ affecting the frequencies of TM/TE modes differently, thereby creating the possibility for mode interaction.
upperbranchupperbranch lowerbranchlowerbranch
[dashed]
$\delta=0$
Refer to caption Refer to caption
[dashed]
$\delta=0.21$
Refer to caption Refer to caption
[dashed]
$\delta=0.23$
Refer to caption Refer to caption
[dashed]
$\delta=0.25$
Refer to caption Refer to caption
Figure 9: Field functions of (2,6)-TM and (2,2)-TE modes for (a,b) $\delta=0$ and (c-h) at the ARC (see Fig. 8(a,b)). The fields are normalized so that $\max|E_{j}|^{2}=1$ and $\max|B_{j}|^{2}=1$, where the subscript $j$ denotes the three field components.
In the following, we concentrate on a specific ARC between a $(1,6)$-TM mode and a $(1,2)$-TE mode, as it is the most common ARC type. Another ARC, where one TM and two TE modes are involved, is shown in the supplementary material in Appendix D. The eigenvalue evolution of the two modes close to the ARC is shown in Fig. 8(a,b). First consider $\mathrm{Re}\,\Omega$ in Fig. 8(a). The two frequency branches approach each other with increasing $\delta$. They are closest at $\delta\approx 0.23$, where both modes exhibit TE polarization. For $\delta<0.23$ the upper (lower) branch is TE (TM) polarized. They switch polarization for $\delta>0.23$. The imaginary component in Fig. 9(b) shows a level crossing. This is expected according to the effective Hamiltonian in Eq. (17). When an (anti-)crossing in (real) imaginary components is present, the coupling is called strong coupling. A second crossing also occurs in the imaginary part at $\delta\approx 0.255$ that is not related to the ARC being considered, but another ARC that is adjacent.
@@
The landscape, i.e., the Riemann surface, close to the EP of the $(1,6)$-TM and $(1,2)$-TE modes is shown in Fig. 8(c,d). The solid black outlines show that the EP needs to be encircled twice in order to return to the starting position. The dashed lines show where $\mathrm{Re}\,(\mathrm{Im}\,)\Delta\Omega=0$, indicating the switch from strong ($\Delta\mathrm{Im}\,\Omega=0$) to weak coupling ($\Delta\mathrm{Re}\,\Omega=0$) when decreasing the refractive index. [72, 76] In the strong coupling regime the parameters in Eq. (17) are such that $2\sqrt{|VW|}>|\mathrm{Im}\,(\Omega^{(1)}-\Omega^{(2)})|$, which results in the level crossing occurring in $\mathrm{Re}\,\Omega$. [72, 76] This is the case for $n>2.4335$ in (see Figs. 8(a,c)) for the considered TM-TE mode pair. The weak coupling regime with $2\sqrt{|VW|}<|\mathrm{Im}\,(\Omega^{(1)}-\Omega^{(2)})|$ is present for $n<2.4335$ (see Fig. 8(d)).
The polarization changes are visible in the wave functions in Fig. 9. The upper branch for $\delta=0$ in Fig. 9(a) is distinctly TE polarized, as the out-of-plane component of the magnetic field dominates the mode structure. The lower branch is TM polarized, see Fig. 9(b). When $\delta$ is increased, the change in mode morphology takes place as discussed in Sec. III.4. The intensity distribution of the TE mode shifts to lower $z$ where the radius increases due to its low $\eta$ (see Fig. 9(c)). The TM mode has an intensity maximum close to the top due to its high $\eta$ (see Fig. 9(d)). Near the ARC in Figs. 9(e,f) both $E_{\parallel}$ and $B_{\parallel}$ are of significant intensity. The TM and TE modes hybridize, which causes both TM and TE components being present. For larger deformations (see Figs. 9(g,h)) the polarization of the two branches switches.
It is evident that EPs govern the coupling between TM and TE modes. The mode pairs whose crossings do not show ARC behavior in Fig. 7 could perhaps be coupled if an appropriate deformation were chosen. Their coupling is also governed by EPs, even though their signatures are not visible here. Note that the ARC shown here is just one example. Numerous other avoided crossings are occurring in the examined parameter space. This includes triplet-mode interaction, an example of which is shown in Appendix D. The underlying EP of third-order could be accessed by varying a third parameter, such as the cavity height, which has proven a useful parameter to control mode coupling in acoustic cavities [77].
@@
upperbranchupperbranch lowerbranchlowerbranch
[dashed]
$\delta=0$
Refer to caption Refer to caption
[dashed]
$\delta=0.21$
Refer to caption Refer to caption
[dashed]
$\delta=0.23$
Refer to caption Refer to caption
[dashed]
$\delta=0.25$
Refer to caption Refer to caption
Figure 10: Husimi functions of modes corresponding to ARC (see Figs. 8(a,b)). The red (blue) figures correspond to the $E_{\parallel}$ ($B_{\parallel}$) field component used for calculation The field functions are shown in Fig. 9, where the numerals in curly brackets refer to the respective parameter.
The Husimi function signatures of the coupled modes are shown in Fig. 10. We choose the Husimi function at $\bm{q}/R_{0}=(0,2.5)$, as nearly all wave functions have a nonzero intensity in the vicinity. The red (blue) colors show the Husimi functions calculated using $E_{\parallel}$ ($B_{\parallel}$). The signatures for the dominant field component of the uncoupled modes at $\delta=0$ are as expected from Fig. 6.
The dominant out-of-plane component of the TE mode in Fig. 9(a) is the $B_{\parallel}$ field. Its Husimi function signature (see Fig. 10(a)) consists of two intensity maxima, which merge to form one broad intensity maximum. This corresponds to the behavior observed in Fig. 6(c) due to the low $\eta$ value. The low-intensity $E_{\parallel}$ field component consists of small fluctuations due to the TE polarization of the mode, which results in a Husimi function structure resembling a resonance with high $\eta$ (compare to Fig. 5(c)).
The uneven spacing of the intensity maxima in $E_{\parallel}$ leads to broader peaks in the Husimi function. The phase space of the TM polarized mode in Fig. 9(b) (left panel) shows the typical behavior of $\eta=6$ in $E_{\parallel}$, as it corresponds to the behavior in Fig. 5(c). Due to the small intensity fluctuations in the $B_{\parallel}$ component (right panel) the Husimi function yields four intensity maxima.
In Figs. 10(c,d), the phase space signatures and the mode morphologies change according to Fig. 6. Furthermore, both the $E_{\parallel}$ and $B_{\parallel}$ components become significant near the ARC. This is seen, for example, in Figs. 9(e,f) and 10(e,f): The electric fields and phase space representations of both eigenmodes display TM mode characteristics, while the magnetic field (and its Husimi function) show TE behavior. The mode hybridization caused by the coupling is thus expressed in real and in phase space. When the system moves away from the ARC, the polarization of the modes increases. The Husimi function signatures in Figs. 10(g,h) reflect the behavior in Figs. 10(c,d). Now the polarization is switched, which is reflected in phase space. For fields and Husimi functions at the EP we refer to Appendix E.